# Frequency-Explicit Shape Holomorphy in Uncertainty Quantification for Acoustic Scattering

### v1

@@ 5 FE Discretization: $k$-Explicit Error Bounds
$\int_{\Omega_{\rm tr}}\bigg{(}k^{-2}\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}-\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}\bigg{)}{\rm d}\widehat{{\boldsymbol{x}}}=\widehat{L}(\widehat{v})$ (5.1)
@@
Here we have followed, e.g., [4, §2], [31, §1.2] and, starting from $\widetilde{\sigma}$, defined $\sigma$ in terms of $\widetilde{\sigma}$. Alternatively, one can start from a non-decreasing function $\sigma$ and define $\widetilde{\sigma}$ such that the first equation in (5.4) holds; see, e.g., [20, §3], [48, §2], [42, §4], [12, §2]. The notation $\alpha$ and $\beta$ is also used by [48, 49].
@@ 5.2 Properties of the sesquilinear form and solution operator of the PML problem
This follows by combining Lemma 4.4 and the analogous result about the sign-definiteness of the PML coefficient, see, e.g., [32, Lemma 2.3] (where $\widetilde{\sigma}(r)=f_{\theta}(r)/r$ in the notation of [32, Lemma 2.3]). In using Lemma 4.4, we use the fact that the constants $C_{\operatorname{Im}},k_{0}$ from Corollary 4.2 are such that if $\mathfrak{r}\in\mathcal{A}$ defined by (4.7), then the assumptions of Lemma 4.4 are satisfied. ∎
@@
Given $n_{i}<1,$ $d=2$ or $3$, and $C_{\operatorname{Re}}>0$, let ${\mathcal{A}}$ and $C_{\operatorname{Im}}$ be as in (4.7), Corollary 4.2. Then, given $\widetilde{\sigma}\in C^{3}(0,\infty)$ as in (5.3), there exists $C,k_{1}\geq\frac{1}{2}$ such that, for all $\mathfrak{r}\in{\mathcal{A}}$, for any $\widehat{L}\in(H^{1}_{k}(B_{2}))^{*}$, and for all $k\geq k_{1}$, the solution $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ of (5.1) exists, is unique, and satisfies the bound
@@
This follows from [31, Theorem 1.6], where $k_{1}$ is taken as the maximum of $k_{0}$ from Corollary 4.2 and the corresponding lower bound on $k$ coming from [31, Theorem 1.6].
Indeed, [31, Theorem 1.6] proves that the solution operator of the PML problem is bounded by the solution operator of the original (i.e., non-truncated) problem for all scattering problems in the black-box scattering framework of [60].
The variational problem (4.2) does not fit in the black-box framework as described in, e.g., [28, Chapter 4], since the PDE is not formally self-adjoint when $\operatorname{Im}\mathfrak{r}\neq 0$ (because the coefficients are complex). Nevertheless, the results from the black-box framework required to prove [31, Theorem 1.6] still hold. Indeed, [31, Theorem 1.6] follows from [31, Lemma 3.3], and the two ingredients of this lemma that depend on the contents of the black box are (i) that the PML problem is Fredholm of index zero, and (ii) agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems (see, e.g., [28, Theorem 4.37]).
Regarding (i): for the variational problem (5.1), this property follows from Corollary 5.3 and, e.g., [50, Theorem 2.34]. Regarding (ii): inspecting the proof of [28, Theorem 4.37], we see that the only place where this proof uses that the contents of the black-box are self adjoint is in finding a complex $k$ where the (unscaled) problem can be shown to have a unique solution and satisfies the natural bound on the solution operator (in terms of $|k|$-dependence); see [28, Lemma 4.3]. In our case, if $k={\rm i}\lambda$, then one can prove directly from the variational formulation (using Lemma 5.2 and the definition of $\widehat{n}$ (2.9b)) that the problem (4.2) with $\mathfrak{r}\in{\mathcal{A}}$ has a unique solution and satisfies the bound (4.5) with $k$ replaced by $\lambda$. ∎
@@
References for the proof.
This result is a special case of [31, Theorem 1.4]. Just like [31, Theorem 1.6] (which we used in the proof of Theorem 5.4) the proof of [31, Theorem 1.4] uses agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems [28, Theorem 4.37]; this result holds for our problem as explained in the proof of Theorem 5.4. ∎
@@ 5.4 The accuracy of the $h$-FEM approximation of the PML solution
Assumption 5.7 holds when $(V_{h})_{h>0}$ consists of functions that are continuous in $\Omega_{\rm tr}$ and piecewise polynomials of total degree $p$ on a shape-regular family of simplicial triangulations of $\Omega_{\rm tr}$, indexed by the meshwidth $h$; see, e.g., [18, Theorem 17.1], [5, Proposition 3.3.17].
@@
Our result about $k$-explicit convergence of the $h$-FEM (Theorem 5.8 below) requires that, for $\gamma\geq p-1$, $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)\in C^{\gamma,1}$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\in C^{\gamma-1,1}$ with the corresponding norms bounded independently of $\mathfrak{r}$ (the reason for this is explained in the discussion preceding Lemma 5.10 and the proof of Theorem 5.8). For simplicity, we assume that $\gamma=p-1$, i.e., given the regularity of $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$, we take the minimal polynomial degree to obtain the best-possible FEM convergence result for that regularity.
@@
These thresholds are observed empirically to be sharp (going back to the work of Ihlenburg and Babuška in 1-d [43, 44]); see the discussion in [33, §1.3] and the references therein.
@@
A key ingredient in the proofs of Theorem 5.8 is the standard elliptic regularity shift property for the highest-order terms in the PDE. We record this result for our problem (i.e., the PML problem (5.1)), emphasizing that the constants in the bound are independent of $\mathfrak{r}$ for $\mathfrak{r}\in{\mathcal{A}}_{p}$.
@@
By, e.g., [50, Theorem 4.18] the bound (5.18) holds if (i) (5.5) holds, (ii) $\partial\Omega_{\rm tr}$ is $C^{p,1}$, and (iii) $\widehat{{\mathbf{A}}}_{\rm PML}\in C^{p-1,1}(\Omega_{\rm tr})$ with norm bounded independently of $\mathfrak{r}$. The property (i) holds by Lemma 5.2 and the property (ii) holds by assumption. For the property (iii), the fact that $\widetilde{\sigma}\in C^{p,1}$ implies that $\widehat{{\mathbf{A}}}_{\rm PML}$ (defined by (5.2)) restricted to the PML region $|\widehat{{\boldsymbol{x}}}|\geq R_{1}$ is $C^{p-1,1}$. Arguing as in Lemma 4.5, we have that if $\mathfrak{r}\in\mathcal{A}_{p}$ then $\widehat{{\mathbf{A}}}_{\rm PML}$ restricted to $|\widehat{{\boldsymbol{x}}}|\leq R_{1}$ is in $C^{p,1}$ with norm bounded independently of $\mathfrak{r}$. ∎
@@
Corollary 5.3 immediately shows that [33, Equations 1.6 and 1.7] are satisfied (i.e., the sesquilinear form is continuous and satisfies a Gårding inequality). For the elliptic regularity assumptions of [33, Assumptions 1.2 and 1.6], we need that there exists $C>0$ such that, for $j=1,\ldots,p+1$ and $\mathfrak{r}\in{\mathcal{A}}_{p}$,
@@
Now, since $\widehat{n}_{\rm PML}\in C^{p-1}$ with norm bounded independently of $\mathfrak{r}\in{\mathcal{A}}_{p}$, by, e.g., [37, Theorem 1.4.1.1, page 21] there exists $C>0$ such that
@@
Assumption 5.7 is [33, Assumption 4.8], and then (5.12)-(5.15) follow from [33, Equations 4.16, 4.17, and 4.19]. The bound (5.16) is not stated explicitly in [33], but follows from the displayed equation before [33, Remark 2.3] (one repeats the arguments that obtain (5.15) from (5.12), but now one starts from (5.13)). ∎
@@ 6.1 Definition of and expressions for the far-field pattern
(this limit exists and is a smooth function of ${\boldsymbol{q}}$ by, e.g., [21, Corollary 3.7]).
@@
We now give an expression for the far-field pattern of an outgoing Helmholtz solution from [54, Theorem 2.2] (written in a slightly-more general way here).
@@
Suppose that, for some $R_{0}>0$, $v\in C^{2}(\mathbb{R}^{d}\setminus B_{R_{0}})$ satisfies the Helmholtz equation $(-k^{-2}\Delta-1)v=0$ in $\mathbb{R}^{d}\setminus\overline{B_{R_{0}}}$ and the Sommerfeld radiation condition (2.6). Let $\psi\in C^{2}(\mathbb{R}^{d};[0,1])$ be such that $\psi\equiv 0$ in a neighbourhood of $B_{R_{0}}$ and $\psi\equiv 1$ on $(B_{R_{1}})^{c}$ for some $R_{1}>R_{0}$. Then
$v_{\infty}({\boldsymbol{q}})=C(d,k)\int_{\operatorname{supp}\nabla\psi}v({\boldsymbol{y}})\Big{(}\Delta\psi({\boldsymbol{y}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi({\boldsymbol{y}})\Big{)}\exp\big{(}-{\rm i}k{\boldsymbol{y}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}{\boldsymbol{y}},$ (6.2)
where
@@
where $\nu({\boldsymbol{y}}):={\boldsymbol{y}}/|{\boldsymbol{y}}|$; see, e.g., [21, Theorem 3.1]. Let $w=v\psi$ and observe that $w\in C^{2}(\overline{B_{R}})$ since $\psi\equiv 0$ on a neighbourhood of $B_{R_{0}}$. With this choice of $w$, as $R\to\infty$, the integral over $\partial B_{R}$ tends to zero, since both $v$ and $G({\boldsymbol{x}},\cdot)$ (for ${\boldsymbol{x}}$ fixed) satisfy the Sommerfeld radiation condition (see, e.g., [21, Last equation in the proof of Theorem 3.3]). Now,
@@
Letting $|{\boldsymbol{x}}|\to\infty$ and using (6.1) and (when $d=2$) the large-argument asymptotics of $H_{0}^{(1)}(\cdot)$ (see, e.g., [56, Equation 10.17.5]), we find that

### v2

@@ 5 Recap of $k$-explicit Finite-Element Error Bounds
$\boxed{\int_{\Omega_{\rm tr}}\bigg{(}k^{-2}\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}-\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}\bigg{)}{\rm d}\widehat{{\boldsymbol{x}}}=\widehat{L}(\widehat{v})}$ (5.1)
@@
Here we have followed, e.g., [4, §2], [33, §1.2] and, starting from $\widetilde{\sigma}$, defined $\sigma$ in terms of $\widetilde{\sigma}$. Alternatively, one can start from a non-decreasing function $\sigma$ and define $\widetilde{\sigma}$ such that the first equation in (5.4) holds; see, e.g., [21, §3], [51, §2], [44, §4], [13, §2]. The notation $\alpha$ and $\beta$ is also used by [51, 52].
@@ 5.2 Properties of the sesquilinear form and solution operator of the PML problem
The sign property (5.5) with the constant $C$ not uniform in $\mathfrak{r}\in\mathcal{A}$ is given in, e.g., [34, Lemma 2.3] (where $\widetilde{\sigma}(r)=f_{\theta}(r)/r$ in the notation of [34, Lemma 2.3]). Since $\mathfrak{r}$ does not enter the definition of the PML scaling function $\widetilde{\sigma}$, the fact that $C$ in (5.5) can be taken to be independent of $\mathfrak{r}\in\mathcal{A}$ follows from Lemma 4.4. In using Lemma 4.4, we use the fact that, by design, the constants $C_{\operatorname{Im}},k_{0}$ from Corollary 4.2 are such that if $\mathfrak{r}\in\mathcal{A}$ defined by (4.7), then the assumptions of Lemma 4.4 are satisfied. ∎
@@
Given $n_{i}<1,$ $d=2$ or $3$, and $C_{\operatorname{Re}}>0$, let ${\mathcal{A}}$ and $C_{\operatorname{Im}}$ be as in (4.7). Then, given $\widetilde{\sigma}\in C^{3}(0,\infty)$ as in (5.3), there exists $C,k_{1}\geq\frac{1}{2}$ such that, for all $\mathfrak{r}\in{\mathcal{A}}$, for any $\widehat{L}\in(H^{1}_{k}(B_{2}))^{*}$, and for all $k\geq k_{1}$, the solution $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ of (5.1) exists, is unique, and satisfies the bound
@@
The result with $C$ in (5.6) not uniform in $\mathfrak{r}\in\mathcal{A}$ follows from [33, Theorem 1.6]. Indeed, [33, Theorem 1.6] proves that the solution operator of the PML problem is bounded by the solution operator of the original (i.e., non-truncated) problem for all scattering problems belonging to the black-box scattering framework of [63].
We now justify that $C$ in (5.6) can be taken to be independent of $\mathfrak{r}\in\mathcal{A}$. An apparent difficulty is that the variational problem (4.2) does not fit in the black-box framework as described in, e.g., [30, Chapter 4], since the PDE is not formally self-adjoint when $\operatorname{Im}\mathfrak{r}\neq 0$ (because the coefficients are complex). Nevertheless, the results from the black-box framework required to prove [33, Theorem 1.6] still hold. Indeed, [33, Theorem 1.6] follows from [33, Lemma 3.3], and the two ingredients of this lemma that depend on the contents of the black box are (i) that the PML problem is Fredholm of index zero, and (ii) agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems (see, e.g., [30, Theorem 4.37]).
Regarding (i): for the variational problem (5.1), this property holds since the sesquilinear form is continuous and satisfies a Gårding inequality (by Corollary 5.3) and the solution is unique for each $\mathfrak{r}\in\mathcal{A}$ by the a priori bound of [33, Theorem 1.6] (it does not matter that the constant a priori depends on $\mathfrak{r}\in\mathcal{A}$ to establish this uniqueness).
Regarding (ii): inspecting the proof of [30, Theorem 4.37], we see that the only place where this proof uses that the contents of the black-box are self adjoint is in finding a complex $k$ where the (unscaled) problem can be shown to have a unique solution and satisfies the natural bound on the solution operator (in terms of $|k|$-dependence); see [30, Lemma 4.3] (the existence of such a complex $k$ then allows one to use analytic Fredholm theory; see, e.g., [30, Theorem C.8]). In our case, if $k={\rm i}\lambda$, then one can prove directly from the variational formulation (i.e., by integration by parts), using Lemma 5.2 and the definition of $\widehat{n}$ (2.9b), that the problem (4.2) with $\mathfrak{r}\in{\mathcal{A}}$ has a unique solution and satisfies the bound (4.5) with $k$ replaced by $\lambda$. ∎
@@
Proof.
The result with the constants $C_{\rm PML,1},C_{\rm PML,2},k_{1}$ not uniform in $\mathfrak{r}\in\mathcal{A}$ follows from [33, Theorem 1.4]. The same arguments in the proof of Theorem 5.4 above show that $C_{\rm PML,1},C_{\rm PML,2},k_{1}$ can be taken independent of $\mathfrak{r}\in\mathcal{A}$. Indeed, like [33, Theorem 1.4], [33, Theorem 1.6] uses agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems [30, Theorem 4.37]; this result holds for our problem as explained in the proof of Theorem 5.4. ∎
@@ 5.4 The accuracy of the $h$-FEM approximation of the PML solution
 [^6] Here, and throughout, the notion “$h$-FEM” refers to the situation where p is fixed and accuracy is increased by varying $h$.
@@
Assumption 5.7 holds when $(V_{h})_{h>0}$ consists of functions that are continuous in $\Omega_{\rm tr}$ and piecewise polynomials of total degree $p$ on a shape-regular family of simplicial triangulations of $\Omega_{\rm tr}$, indexed by the meshwidth $h$; see, e.g., [19, Theorem 17.1], [5, Proposition 3.3.17].
@@
Our result about $k$-explicit convergence of the $h$-FEM (Theorem 5.8 below) requires that, for $\gamma\geq p-1$, $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)\in C^{\gamma,1}$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\in C^{\gamma-1,1}$ with the corresponding norms bounded independently of $\mathfrak{r}$ (the reason for this is explained in the discussion preceding Lemma 5.10 and the proof of Theorem 5.8). For simplicity, we assume that $\boxed{p=\gamma+1}$, i.e., given the regularity of $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$, we take the minimal polynomial degree to obtain the best-possible FEM convergence result for that regularity. Since the regularity of $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ depend on the regularity of the radial displacement function (via (2.9)), this means that we are selecting the minimal polynomial degree to obtain the best-possible FEM convergence result for the given regularity of the transmission interface.
@@
These thresholds are observed empirically to be sharp (going back to the work of Ihlenburg and Babuška in 1-d [45, 46]); see the discussion in [35, §1.3] and the references therein.
@@
The result of Theorem 5.8 with the constants $C_{\rm FEM,j},j=1,2,3,4$, depending on $\mathfrak{r}\in\mathcal{A}$ is a direct consequence of the main result of [35]. The assumptions in [35] are that (i) the sesquilinear form is continuous and satisfies a Gårding inequality, and (ii) the highest-order terms in the underlying PDE satisfy the natural elliptic-regularity shift, and the constants in the FEM-error bounds in [35] then only depend on the constants in (i) and (ii). To prove Theorem 5.8 therefore, we need to show that the properties (i) and (ii) hold with constants uniform in $\mathfrak{r}\in\mathcal{A}_{p}$. For (i), this property follows immediately from Corollary 5.3. We therefore now focus on (ii).
@@
By the standard elliptic-regularity shift result (see, e.g., [53, Theorem 4.18]) the bound (5.18) holds if (i) (5.5) holds, (ii) $\partial\Omega_{\rm tr}$ is $C^{p,1}$, and (iii) $\widehat{{\mathbf{A}}}_{\rm PML}\in C^{p-1,1}(\Omega_{\rm tr})$ with norm bounded independently of $\mathfrak{r}$. The property (i) holds by Lemma 5.2 and the property (ii) holds by assumption. For the property (iii), the fact that $\widetilde{\sigma}\in C^{p,1}$ implies that $\widehat{{\mathbf{A}}}_{\rm PML}$ (defined by (5.2)) restricted to the PML region $|\widehat{{\boldsymbol{x}}}|\geq R_{1}$ is $C^{p-1,1}$. Arguing as in Lemma 4.5, we have that if $\mathfrak{r}\in\mathcal{A}_{p}$, then $\widehat{{\mathbf{A}}}_{\rm PML}$ restricted to $|\widehat{{\boldsymbol{x}}}|\leq R_{1}$ is in $C^{p,1}$ with norm bounded independently of $\mathfrak{r}$. ∎
With Corollary 5.3 and Lemma 5.10 in hand, we can now prove Theorem 5.8.
@@
Corollary 5.3 immediately shows that [35, Equations 1.6 and 1.7] are satisfied (i.e., the sesquilinear form is continuous and satisfies a Gårding inequality). For the elliptic regularity assumptions of [35, Assumptions 1.2 and 1.6], we need that there exists $C>0$ such that, for $j=1,\ldots,p+1$ and $\mathfrak{r}\in{\mathcal{A}}_{p}$,
@@
Now, since $\widehat{n}_{\rm PML}\in C^{p-1}$ with norm bounded independently of $\mathfrak{r}\in{\mathcal{A}}_{p}$, by a classic result about the Sobolev norm of a product (see, e.g., [40, Theorem 1.4.1.1, page 21]) there exists $C>0$ such that
@@
Assumption 5.7 is [35, Assumption 4.8], and then (5.12)-(5.15) follow from [35, Equations 4.16, 4.17, and 4.19]. The bound (5.16) is not stated explicitly in [35], but follows from the displayed equation before [35, Remark 2.3] (one repeats the arguments that obtain (5.15) from (5.12), but now one starts from (5.13)). ∎
@@ 6.1 Definition of and expressions for the far-field pattern
(this limit exists and is a smooth function of ${\boldsymbol{q}}$ by, e.g., [22, Corollary 3.7]).
@@
We now give an expression for the far-field pattern of an outgoing Helmholtz solution from [57, Theorem 2.2] (written in a slightly-more general way here).
@@
Suppose that, for some $R_{0}>0$, $v\in C^{2}(\mathbb{R}^{d}\setminus B_{R_{0}})$ satisfies the Helmholtz equation $(-k^{-2}\Delta-1)v=0$ in $\mathbb{R}^{d}\setminus\overline{B_{R_{0}}}$ and the Sommerfeld radiation condition (2.6). Let $\psi\in C^{2}(\mathbb{R}^{d};[0,1])$ be such that $\psi\equiv 0$ in a neighbourhood of $B_{R_{0}}$ and $\psi\equiv 1$ on $(B_{R_{1}})^{c}$ for some $R_{1}>R_{0}$.
Then
$\boxed{v_{\infty}({\boldsymbol{q}})=C(d,k)\int_{\operatorname{supp}\nabla\psi}v({\boldsymbol{y}})\Big{(}\Delta\psi({\boldsymbol{y}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi({\boldsymbol{y}})\Big{)}\exp\big{(}-{\rm i}k{\boldsymbol{y}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}{\boldsymbol{y}}}\;,$ (6.2)
where ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$ and
@@
where $\nu({\boldsymbol{y}}):={\boldsymbol{y}}/|{\boldsymbol{y}}|$; see, e.g., [22, Theorem 3.1]. Let $w=v\psi$ and observe that $w\in C^{2}(\overline{B_{R}})$ since $\psi\equiv 0$ on a neighbourhood of $B_{R_{0}}$. With this choice of $w$, as $R\to\infty$, the integral over $\partial B_{R}$ tends to zero, since both $v$ and $G({\boldsymbol{x}},\cdot)$ (for ${\boldsymbol{x}}$ fixed) satisfy the Sommerfeld radiation condition (see, e.g., [22, Last equation in the proof of Theorem 3.3]). Now,
@@
Letting $|{\boldsymbol{x}}|\to\infty$ and using (6.1) and (when $d=2$) the large-argument asymptotics of $H_{0}^{(1)}(\cdot)$ (see, e.g., [59, Equation 10.17.5]), we find that